A Complete Bibliography of ACM Transactions on Speech and Language Processing (TSLP)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 April 2024 Version 1.32

Title word cross-reference

N [GKVS08].

-gram [GKVS08].

ABCs [BKTN11]. acceptability [DB13]. accuracy [EBR11]. accuracy-enhanced [EBR11]. Acoustic [MMS06]. across [CLS13b, CHK^{+}07, SL06]. Active [ZF12, HTRT06, Zha13, ZWHM10, ZM12]. actor [JTY11]. Adaptive [RSM07]. adaptivity [LP11]. affect [Zha13]. affect-sensing [Zha13]. Age [MT11]. aided [WS06]. algorithm [JTY11]. algorithms [OBC13]. Analysis [RVKM12, YBKR07]. annotation [ZF12, ZWHM10]. applications [GFL06, PGNB11]. approach [HTRT06]. Arabic [EBR11]. aspect [ZLG13]. aspect-guided [ZLG13]. assessing [BKTN11, DMS11]. assessment [BCW^{+}11]. audio [TK11]. audit [MKP^{+}09]. Author [Van07]. Automatic [ALSZ06, KR05, GLR07, MHS^{+}11, ZF12]. Automatically [BKTN11]. aware [CD11].

based [AS12, CLZ13, CHK^{+}07, GY11, LK05, NZ13, UPN10, Zha13, ZWHM10, ZM12]. batch [PGCFB11]. behavior [RE12]. belief [JTY11]. benchmark [OBC13]. bigram [TZ07]. bilingual [MDTK10, PP11]. BLEU [SWY13]. block [TZ07]. bootstrapping [GSD09]. Brains [MDTK10]. brawn [MDTK10]. Broad [SL06]. canonicalization [RZ13]. case [BM13, KBM13]. categorization [ALSZ06, GSD09]. challenges [BACMW12]. child [PGNB11, WSB^{+}11]. child-machine [PGNB11]. child-robot [WSB^{+}11]. children [BKTN11, BCW^{+}11, PGNB11, WSB^{+}11]. Chinese [YBKR07, ZYS08]. choice [Ink07]. classification [MMS06, YBKR07, ZM12]. Classifying [GBM11]. clustering [CLZ13]. co [CLZ13]. co-clustering [CLZ13]. Cognitive [RZ13]. coherence [SKH^{+}08, ZLG13]. collections [DYIN13]. collocations [LBN13]. Combining [CLZ13, CLS13a]. communication [SPF^{+}12, THT^{+}05]. communicator [HOWM07]. comparable [MDTK10]. comparative [UPN10]. Comparing [AL11]. Composition [BM13]. compound [CLS13a]. compounds [KB13, KBM13, NH13]. comprehensive [UPN10]. computational [SKTK13]. computer [WS06]. computer-aided [WS06]. conditional [CLS13a]. Confidence [GFL06, ZWHM10]. Confidence-based [ZWHM10]. Construction [YBKR07]. constructions [DB13, VTZ13]. content [NPM07, ZLG13]. content-level [ZLG13]. context [AS12]. Contextual [Zha13]. control [CD11, MMHD13]. conversational [BOS^{+}07, WCB^{+}11]. corpora [MDTK10, SKH^{+}08]. correction [RSM07]. Cost [NR12]. coverage [SL06]. crawled [RSM07]. criteria [ZWHM10]. critic [JTY11].

DASH [MT11]. data [ZWHM10]. dataset [OBC13]. decision [MMHD13, MSK^{+}11, MKP^{+}09]. decoding [WSB^{+}11]. dependency [YBKR07]. dependent [RSM07]. detect [VTZ13]. detection [CLZ13, MT11, WSB^{+}11]. dialog [HOWM07, RE12]. dialogue [AL11, CD11, GY11, GBM11, HMNA04, HWP07, JTY11, LP11, MMHD13, MSK^{+}11, PGCFB11]. dictionaries [HWP07, RSM07]. difficulty [CLS13b]. dimensional [GBM11]. disambiguation [AS12]. discourse [HMNA04]. disorder [MHS^{+}11]. disparity [DYIN13]. Distributed [SPF^{+}12]. do [Chu13]. document [DYIN13]. domain [ALSZ06, RSM07]. domain-dependent [RSM07]. domain-specific [ALSZ06]. domains [SL06]. drama [Zha13].

Editorial [The13]. educational [HM11]. Effective [GY11]. efficient [PGCFB11, THT^{+}05]. elementary [WCB^{+}11]. empirical [SKH^{+}08]. English [DB13, SKH^{+}08, VTZ13]. enhanced [EBR11]. entity [GLR07]. environment [MMS06]. essays [KBM13]. estimation [GFL06, ZM12]. evaluate [CLS13b]. Evaluating [HMNA04]. evaluation [GKVS08, HOWM07, MKP^{+}09, NPM07, UPN10]. expansion [ALSZ06]. exploration [Van07]. expressions [Chu13, KBM13, NZ13, RVK13]. extension [HOWM07]. extraction [GLR07, KVM11]. extractive [ZF12]. Extrinsic [MKP^{+}09].

factors [DB13]. features [KR05]. fields [CLS13a]. Finnish [TK11]. FLORA [BCW^{+}11]. flow [FN06]. Fluent [BCW^{+}11]. fonts [VK05]. football [BACMW12]. framework [BM13]. Functions [NR12].

game [SKH^{+}08]. game-theoretic

[SKH^{+}08]. gender [MT11]. generation [BACMW12, HWP07]. graffiti [THT^{+}05]. gram [GKVS08]. graphs [GKVS08]. guided [ZLG13].

handling [GY11]. hidden [GY11, FN06]. hierarchical [CD11]. High [WS06, GBM11]. high-dimensional [GBM11]. High-quality [WS06]. human [NPM07, THT^{+}05]. Hungarian [VTZ13].

I-DASH [MT11]. identify [HM11]. importance [AS12]. Improving [GSD09]. improvisation [Zha13]. Incorporating [NPM07]. individuality [VK05]. influence [GLR07]. information [CLS13b, DYIN13, GY11]. information-theoretic [CLS13b]. instability [ZM12]. interaction [PGNB11, WSB^{+}11]. internal [NZ13]. interpretation [NH13]. Introduction [LP11, PGNB11, RVK13]. Investigating [RVKM12]. issue [LP11, PGNB11, RVK13].

Japanese [SKH^{+}08].

keyword [WSB^{+}11]. know [Chu13].

labeling [KVM11]. Language [RVKM12, BOS^{+}07, HTRT06, KVM11, LK05, PGNB11, RZ13, WS06]. languages [CHK^{+}07, SL06]. large [SKH^{+}08]. Lattice [SWY13]. lattices [CLS13a]. Learning [HM11, MMHD13, VTZ13, AL11, CL07, CD11, GSD09, HWP07, JTY11, LP11, PGCFB11, WS06, ZF12, Zha13, ZWHM10, ZM12]. learning-based [Zha13]. letter [BKTN11]. letter-names [BKTN11]. letter-sounds [BKTN11]. level [ZLG13]. Lexical [DB13, KR05]. lexicons [ALSZ06, PP11]. light [EBR11, VTZ13]. like [TK11]. linguistic [Van07]. listening [MMHD13]. listening-oriented [MMHD13]. logical [SKTK13].

machine [LP11, PGNB11, SWY13, TZ07, THT^{+}05, ZYS08]. management [PGCFB11]. manager [GY11, HOWM07]. Managing [DYIN13]. many [Chu13]. Markov [FN06, MMHD13]. match [BACMW12]. materials [HM11]. measure [CLS13b]. messages [KR05]. Metaphorical [RVKM12]. method [HWP07, NZ13, NPM07]. methodology [OBC13]. methods [DMS11, UPN10]. metonymy [SKTK13]. mining [HWP07, MDTK10]. model [FB05, Ink07, SKH^{+}08, SKTK13, TZ07]. Modeling [MSK^{+}11, NZ13, BOS^{+}07, CHK^{+}07]. modelled [JTY11]. models [CL07, LK05, LBN13, FN06]. Morph [CHK^{+}07]. Morph-based [CHK^{+}07]. morpheme [CL07, TK11]. morpheme-like [TK11]. morphology [CL07]. multidocument [FN06]. multilingual [DYIN13, FN06, SPF^{+}12]. multimedia [WCB^{+}11]. multiparty [SPF^{+}12]. multiword [Chu13, KBM13, NZ13, RVK13]. My [WCB^{+}11].

named [GLR07]. named-entity [GLR07]. names [BKTN11]. Natural [JTY11, KVM11, LK05, RZ13]. near [Ink07]. near-synonym [Ink07]. NLP [GFL06]. noise [CLZ13]. nominalization [DB13]. noun [KB13, KBM13, NH13].

observable [MMHD13]. old [BACMW12]. One [FN06]. optimization [MSK^{+}11, PGCFB11]. Optimizing [RE12]. oracles [SWY13]. oral [BCW^{+}11, DMS11]. oriented [BACMW12, MMHD13, RE12]. other [NH13]. out-of-vocabulary [CHK^{+}07].

paragraph [SL06]. parameter [Van07]. parameters [JTY11]. paraphrases [NH13]. parsing [CLS13a, CLS13b]. partially [MMHD13]. pattern [NZ13]. pattern-based [NZ13]. PCFG [CLS13a]. people [Chu13]. Perspective [BACMW12, RVKM12]. Perspective-oriented [BACMW12]. phrase [FB05]. POMDP [GY11]. POMDP-based [GY11]. POMDPs [JTY11]. practice [RVK13]. prediction [TZ07]. processes [MMHD13]. processing [HTRT06, LK05, PGNB11]. production [OBC13]. profiles [KBM13]. profiling [Van07]. project [MT11, THT^{+}05]. Promoting [HOWM07]. prosodic [KR05]. prosody [DMS11]. Pyramid [NPM07].

quality [WS06]. queen [HOWM07]. queries [RZ13].

random [CLS13a]. reading [BCW^{+}11, DMS11, MHS^{+}11]. recognition [BOS^{+}07, CLS13a, CHK^{+}07, GLR07, TK11]. referential [SKH^{+}08]. Reinforcement [JTY11, CD11, PGCFB11]. Relation [GLR07]. relations [BM13, KB13, KVM11, PP11]. representation [VK05]. resources [BOS^{+}07]. retrieval [TK11]. reuse [HOWM07]. Reviewers [Ano13]. reviews [HWP07]. revisited [GKVS08]. robot [WSB^{+}11]. role [KVM11]. RST [UPN10]. RST-based [UPN10].

Sample [PGCFB11]. Sample-efficient [PGCFB11]. scenario [WSB^{+}11]. school [WCB^{+}11]. science [WCB^{+}11]. segmentation [CL07, SL06, TK11, ZYS08]. Selection [NR12, NPM07]. Semantic [NH13, PP11, BM13, DB13, KB13, OBC13, RZ13, YBKR07]. semi [ZF12]. semi-automatic [ZF12]. Sense [RVKM12, AS12, KB13]. Sense-to-Sentiment [RVKM12]. sensing [Zha13]. sentential [AS12]. Sentiment [KBM13, RVKM12]. short [OBC13]. similarity [AS12, OBC13]. similarity-based [AS12]. simulations [AL11]. smart [MDTK10]. sounds [BKTN11]. space [GBM11, Van07]. Spatial [KVM11]. Spatially [CD11]. Spatially-aware [CD11]. special [LP11, PGNB11, RVK13]. Specific [NR12, ALSZ06]. Speech [NR12, TK11, BCW^{+}11, BOS^{+}07, CHK^{+}07, PGNB11, SPF^{+}12, THT^{+}05, WS06, WSB^{+}11, ZF12]. speech-to-speech [WS06]. spoken [BKTN11, HTRT06, HOWM07, HMNA04, HWP07, LP11, MSK^{+}11, RE12]. state [GY11]. statistical [FB05, Ink07, TZ07, TK11, ZYS08]. stemmer [EBR11]. stopping [ZWHM10]. story [FN06]. strata [RZ13]. strategy [AL11, MSK^{+}11]. study [BM13]. summaries [BACMW12]. Summarization [GKVS08, CLZ13, FN06, KR05, MKP^{+}09, NPM07, UPN10, ZF12, ZLG13]. support [DB13, MSK^{+}11]. support-verb-nominalization [DB13]. Syllable [NR12]. synonym [Ink07]. Synthesis [NR12]. system [GKVS08]. systems [HMNA04, HWP07, JTY11, LP11, RE12].

taker [KBM13]. taking [RE12]. Tandem [WSB^{+}11]. task [MKP^{+}09, RE12]. task-oriented [RE12]. tasks [BACMW12]. technologies [SPF^{+}12]. term [ALSZ06]. terminology [MDTK10]. test [KBM13, MHS^{+}11]. test-taker [KBM13]. Text [NR12, EBR11, GSD09, OBC13, RSM07, ZM12]. Text-to-Speech [NR12]. theme [CLZ13]. theme-based [CLZ13]. theoretic [CLS13b, SKH^{+}08]. Theoretical [BM13]. theory [RVK13]. topic [LBN13]. transformation [VK05]. translation [FB05, SPF^{+}12, SWY13, TZ07, WS06, ZYS08]. treebank [YBKR07]. treebanks [CLS13b]. turn [RE12]. turn-taking [RE12]. tutor [WCB^{+}11]. Two [DMS11].

Uncertainty [ZM12]. Uncertainty-based [ZM12]. understanding [HMNA04]. Unit [NR12]. units [TK11]. unsegmented [TK11]. Unsupervised [AS12, CL07, GSD09, HWP07]. use [MDTK10, RVK13, YBKR07]. user [AL11, HWP07]. using [AS12, CD11, KR05, MMHD13, NH13, RZ13, SKH^{+}08, TK11]. variability [NZ13]. variation [NPM07]. vectors [AS12]. verb [DB13, VTZ13]. verbal [NH13]. Verification [BKTN11, SKH^{+}08, Van07]. version [MHS^{+}11]. via [GSD09]. virtual [WCB^{+}11, Zha13]. vocabulary [CHK^{+}07]. Voice [VK05]. voicemail [KR05].

Web [BOS^{+}07, LK05, RSM07]. Web-based [LK05]. Web-crawled [RSM07]. Word [KB13, AS12, CLS13a, FB05, ZYS08]. wordto-phrase [FB05]. words [CHK^{+}07].

References

Ai:2011:CUS

[AL11] Hua Ai and Diane Litman.

Comparing user simulations for dialogue strategy learning. ACM Transactions on Speech and Language Processing (TSLP), 7(3):9:1–9:??, May 2011. CODEN ???? ISSN 1550-4875.

Avancini:2006:AED

[ALSZ06] Henri Avancini, Alberto Lavelli, Fabrizio Sebastiani, and Roberto Zanoli. Automatic expansion of domain-specific lexicons by term categorization. ACM Transactions on Speech and Language Processing (TSLP), 3 (1):1–30, May 2006. CODEN ???? ISSN 1550-4875.

Anonymous:2013:R

[Ano13] Anonymous. Reviewers. ACM Transactions on Speech and Language Processing (TSLP), 10(2):8:1–8:??, June 2013. CO-DEN ???? ISSN 1550-4875.

Abdalgader:2012:USB

[AS12] Khaled Abdalgader and Andrew Skabar. Unsupervised similarity-based word sense disambiguation using context vectors and sentential word importance. ACM Transactions on Speech and Language Processing (TSLP), 9(1):2:1–2:??, May 2012. CODEN ???? ISSN 1550-4875.

Bouayad-Agha:2012:POG

[BACMW12] Nadjet Bouayad-Agha, Gerard Casamayor, Simon Mille, and Leo Wanner. Perspectiveoriented generation of football match summaries: old tasks, new challenges. ACM Transactions on Speech and Language Processing (TSLP), 9(2): 3:1–3:??, July 2012. CODEN ???? ISSN 1550-4875.

Bolanos:2011:FFO

[BCW^{+}11] Daniel Bolaños, Ronald A. Cole, Wayne Ward, Eric Borts, and Edward Svirsky. FLORA: Fluent oral reading assessment of children’s speech. ACM Transactions on Speech and Language Processing (TSLP), 7(4):16:1– 16:??, August 2011. CODEN ???? ISSN 1550-4875.

Black:2011:AAA

[BKTN11] Matthew P. Black, Abe Kazemzadeh, Joseph Tepperman, and Shrikanth S. Narayanan. Automatically assessing the ABCs: Verification of children’s spoken letternames and letter-sounds. ACM Transactions on Speech and Language Processing (TSLP), 7(4):15:1–15:??, August 2011. CODEN ???? ISSN 1550-4875.

Blanco:2013:CSR

[BM13] Eduardo Blanco and Dan Moldovan. Composition of semantic relations: Theoretical framework and case study. ACM Transactions on Speech and Language Processing (TSLP), 10(4):17:1–17:??, December 2013. CODEN ???? ISSN 1550-4875.

Bulyko:2007:WRL

[BOS^{+}07] Ivan Bulyko, Mari Ostendorf, Manhung Siu, Tim Ng, Andreas Stolcke, and Özgür Çetin. Web resources for language modeling in conversational speech recognition. ACM Transactions on Speech and Language Processing (TSLP), 5(1):1:1–1:25, December 2007. CODEN ???? ISSN 1550-4875.

Cuayahuitl:2011:SAD

[CD11] Heriberto Cuayáhuitl and Nina Dethlefs. Spatially-aware dialogue control using hierarchical reinforcement learning. ACM Transactions on Speech and Language Processing (TSLP), 7 (3):5:1–5:??, May 2011. CO-DEN ???? ISSN 1550-4875.

Creutz:2007:MBS

[CHK^{+}07] Mathias Creutz, Teemu Hirsimäki, Mikko Kurimo, Antti Puurula, Janne Pylkkönen, Vesa Siivola, Matti Varjokallio, Ebru Arisoy, Murat Saraçlar, and Andreas Stolcke. Morphbased speech recognition and modeling of out-of-vocabulary words across languages. ACM Transactions on Speech and Language Processing (TSLP), 5(1):3:1–3:29, December 2007. CODEN ???? ISSN 1550-4875.

Church:2013:HMM

[Chu13] Kenneth Church. How many multiword expressions do people know? ACM Transactions on Speech and Language Processing (TSLP), 10(2):4:1–4:??, June 2013. CODEN ???? ISSN 1550-4875.

Creutz:2007:UMM

[CL07] Mathias Creutz and Krista Lagus. Unsupervised models for morpheme segmentation and morphology learning. ACM Transactions on Speech and Language Processing (TSLP), 4 (1):3:1–3:??, January 2007. CO-DEN ???? ISSN 1550-4875.

Constant:2013:CCR

[CLS13a] Matthieu Constant, Joseph Le Roux, and Anthony Sigogne. Combining compound recognition and PCFG–LA parsing with word lattices and conditional random fields. ACM Transactions on Speech and Language Processing (TSLP), 10(3):8:1–8:??, July 2013. CO-DEN ???? ISSN 1550-4875.

Corazza:2013:ITM

[CLS13b] Anna Corazza, Alberto Lavelli, and Giorgio Satta. An information-theoretic measure to evaluate parsing difficulty across treebanks. ACM Transactions on Speech and Language Processing (TSLP), 9(4):7:1– 7:??, January 2013. CODEN ???? ISSN 1550-4875.

Cai:2013:CCC

[CLZ13] Xiaoyan Cai, Wenjie Li, and Renxian Zhang. Combining co-clustering with noise detection for theme-based summarization. ACM Transactions on Speech and Language Processing (TSLP), 10(4):16:1–16:??, December 2013. CODEN ???? ISSN 1550-4875.

Davis:2013:LSF

[DB13] Anthony R. Davis and Leslie Barrett. Lexical semantic factors in the acceptability of English supportverb-nominalization constructions. ACM Transactions on Speech and Language Processing (TSLP), 10(2):5:1–5:??, June 2013. CODEN ???? ISSN 1550-4875.

Duong:2011:TMA

[DMS11] Minh Duong, Jack Mostow, and Sunayana Sitaram. Two methods for assessing oral reading prosody. ACM Transactions on Speech and Language Processing (TSLP), 7(4):14:1–14:??, August 2011. CODEN ???? ISSN 1550-4875.

Duh:2013:MID

[DYIN13] Kevin Duh, Ching-Man Au Yeung, Tomoharu Iwata, and Masaaki Nagata. Managing information disparity in multilingual document collections. ACM Transactions on Speech and Language Processing (TSLP), 10(1):1:1–1:??, March 2013. CODEN ???? ISSN 1550-4875.

El-Beltagy:2011:AEL

[EBR11] Samhaa R. El-Beltagy and Ahmed Rafea. An accuracyenhanced light stemmer for Arabic text. ACM Transactions on Speech and Language Processing (TSLP), 7(2):2:1–2:??, February 2011. CODEN ???? ISSN 1550-4875.

Federico:2005:WPS

[FB05] Marcello Federico and Nicola Bertoldi. A word-to-phrase statistical translation model. ACM Transactions on Speech and Language Processing (TSLP), 2 (2):1–24, December 2005. CO-DEN ???? ISSN 1550-4875.

Fung:2006:OSO

[FN06] Pascale Fung and Grace Ngai. One story, one flow: Hidden Markov Story Models for multilingual multidocument summarization. ACM Transactions on Speech and Language Processing (TSLP), 3(2):1–16, July 2006. CODEN ???? ISSN 1550-4875.

Gonzalez-Brenes:2011:CDH

[GBM11] José P. González-Brenes and Jack Mostow. Classifying dialogue in high-dimensional space. ACM Transactions on Speech and Language Processing (TSLP), 7(3):8:1–8:??, May 2011. CODEN ???? ISSN 1550-4875.

Gandrabur:2006:CEN

[GFL06] Simona Gandrabur, George Foster, and Guy Lapalme. Confidence estimation for NLP applications. ACM Transactions on Speech and Language Processing (TSLP), 3(3):1–29, October 2006. CODEN ???? ISSN 1550-4875.

Giannakopoulos:2008:SSE

[GKVS08] George Giannakopoulos, Vangelis Karkaletsis, George Vouros, and Panagiotis Stamatopoulos. Summarization system evaluation revisited: N-gram graphs. ACM Transactions on Speech and Language Processing (TSLP), 5(3):5:1–5:??, October 2008. CODEN ???? ISSN 1550-4875.

Giuliano:2007:REI

[GLR07] Claudio Giuliano, Alberto Lavelli, and Lorenza Romano. Relation extraction and the influence of automatic namedentity recognition. ACM Transactions on Speech and Language Processing (TSLP), 5(1):2:1– 2:26, December 2007. CODEN ???? ISSN 1550-4875.

Gliozzo:2009:ITC

[GSD09] Alfio Gliozzo, Carlo Strapparava, and Ido Dagan. Improving text categorization bootstrapping via unsupervised learning. ACM Transactions on Speech and Language Processing (TSLP), 6(1):1:1–1:??, October 2009. CODEN ???? ISSN 1550-4875.

Gavsic:2011:EHD

[GY11] Milica Gavšić and Steve Young. Effective handling of dialogue state in the hidden information state POMDP-based dialogue manager. ACM Transactions on Speech and Language Processing (TSLP), 7(3):4:1–4:??, May 2011. CODEN ???? ISSN 1550-4875.

Hassan:2011:LIE

[HM11] Samer Hassan and Rada Mihalcea. Learning to identify educational materials. ACM Transactions on Speech and Language Processing (TSLP), 8(2): 2:1–2:??, November 2011. CO-DEN ???? ISSN 1550-4875.

Higashinaka:2004:EDU

[HMNA04] Ryuichiro Higashinaka, Noboru Miyazaki, Mikio Nakano, and Kiyoaki Aikawa. Evaluating discourse understanding in spoken dialogue systems. ACM Transactions on Speech and Language Processing (TSLP), 1(1):1–20, November 2004. CODEN ???? ISSN 1550-4875.

Hanna:2007:PER

[HOWM07] Philip Hanna, Ian O’Neill, Craig Wootton, and Michael Mctear. Promoting extension and reuse in a spoken dialog manager: An evaluation of the queen’s communicator. ACM Transactions on Speech and Language Processing (TSLP), 4 (3):7:1–7:??, July 2007. CO-DEN ???? ISSN 1550-4875.

Hakkani-Tur:2006:AAS

[HTRT06] Dilek Hakkani-Tür, Giuseppe Riccardi, and Gokhan Tur. An active approach to spoken language processing. ACM Transactions on Speech and Language Processing (TSLP), 3(3):1–31, October 2006. CODEN ???? ISSN 1550-4875.

Higashinaka:2007:UML

[HWP07] Ryuichiro Higashinaka, Marilyn A. Walker, and Rashmi Prasad. An unsupervised method for learning generation dictionaries for spoken dialogue systems by mining user reviews. ACM Transactions on Speech and Language Processing (TSLP), 4(4):8:1–8:??, October 2007. CODEN ???? ISSN 1550-4875.

Inkpen:2007:SMN

[Ink07]  Diana Inkpen.  A statisti-

cal model for near-synonym choice. ACM Transactions on Speech and Language Processing (TSLP), 4(1):2:1–2:??, January 2007. CODEN ???? ISSN 1550-4875.

Jurvcicek:2011:NAB

[JTY11] Filip Jurvčíček, Blaise Thomson, and Steve Young. Natural actor and belief critic: Reinforcement algorithm for learning parameters of dialogue systems modelled as POMDPs. ACM Transactions on Speech and Language Processing (TSLP), 7(3):6:1–6:??, May 2011. CODEN ???? ISSN 1550-4875.

Kim:2013:WSS

[KB13] Su Nam Kim and Timothy Baldwin. Word sense and semantic relations in noun compounds. ACM Transactions on Speech and Language Processing (TSLP), 10(3):9:1–9:??, July 2013. CODEN ???? ISSN 1550-4875.

Klebanov:2013:SPM

[KBM13] Beata Beigman Klebanov, Jill Burstein, and Nitin Madnani. Sentiment profiles of multiword expressions in test-taker essays: The case of noun–noun compounds. ACM Transactions on Speech and Language Processing (TSLP), 10(3):12:1–12:??, July 2013. CODEN ???? ISSN 1550-4875.

Koumpis:2005:ASV

[KR05] Konstantinos Koumpis and Steve Renals. Automatic summarization of voicemail messages using lexical and prosodic features. ACM Transactions on Speech and Language Processing (TSLP), 2(1):1–24, February 2005. CODEN ???? ISSN 1550-4875.

Kordjamshidi:2011:SRL

[KVM11] Parisa Kordjamshidi, Martijn Van Otterlo, and Marie-Francine Moens. Spatial role labeling: Towards extraction of spatial relations from natural language. ACM Transactions on Speech and Language Processing (TSLP), 8(3):4:1–4:??, December 2011. CODEN ???? ISSN 1550-4875.

Lau:2013:CTM

[LBN13] Jey Han Lau, Timothy Baldwin, and David Newman. On collocations and topic models. ACM Transactions on Speech and Language Processing (TSLP), 10(3):10:1–10:??, July 2013. CODEN ???? ISSN 1550-4875.

Lapata:2005:WBM

[LK05] Mirella Lapata and Frank Keller. Web-based models for natural language processing. ACM Transactions on Speech and Language Processing (TSLP), 2(1):1–31, February 2005. CODEN ???? ISSN 1550-4875.

Lemon:2011:ISI

[LP11] Oliver Lemon and Olivier Pietquin. Introduction to special issue on machine learning for adaptivity in spoken dialogue systems. ACM Transactions on Speech and Language Processing (TSLP), 7(3):3:1– 3:??, May 2011. CODEN ???? ISSN 1550-4875.

Morin:2010:BBU

[MDTK10] Emmanuel Morin, Béatrice Daille, Koichi Takeuchi, and Kyo Kageura. Brains, not brawn: The use of “smart” comparable corpora in bilingual terminology mining. ACM Transactions on Speech and Language Processing (TSLP), 7(1):1:1– 1:??, August 2010. CODEN ???? ISSN 1550-4875.

Maier:2011:AVR

[MHS^{+}11] Andreas Maier, Flonan Hönig, Stefan Steidl, Elmar Nöth, Stefanie Horndasch, Elisabeth Sauerhöfer, Oliver Kratz, and Gunther Moll. An automatic version of a reading disorder test. ACM Transactions on Speech and Language Processing (TSLP), 7(4):17:1–17:??, August 2011. CODEN ???? ISSN 1550-4875.

Murray:2009:ESE

[MKP^{+}09] Gabriel Murray, Thomas Kleinbauer, Peter Poller, Tilman Becker, Steve Renals, and Jonathan Kilgour. Extrinsic summarization evaluation: A decision audit task. ACM Transactions on Speech and Language Processing (TSLP), 6 (2):2:1–2:??, October 2009. CO-DEN ???? ISSN 1550-4875.

Meguro:2013:LCL

[MMHD13] Toyomi Meguro, Yasuhiro Minami, Ryuichiro Higashinaka, and Kohji Dohsaka. Learning to control listening-oriented dialogue using partially observable Markov decision processes. ACM Transactions on Speech and Language Processing (TSLP), 10(4):15:1–15:??, December 2013. CODEN ???? ISSN 1550-4875.

Ma:2006:AEC

[MMS06] Ling Ma, Ben Milner, and Dan Smith. Acoustic environment classification. ACM Transactions on Speech and Language Processing (TSLP), 3(2):1–22, July 2006. CODEN ???? ISSN 1550-4875.

Misu:2011:MSD

[MSK^{+}11] Teruhisa Misu, Komei Sugiura, Tatsuya Kawahara, Kiyonori Ohtake, Chiori Hori, Hideki Kashioka, Hisashi Kawai, and Satoshi Nakamura. Modeling spoken decision support dialogue and optimization of its dialogue strategy. ACM Transactions on Speech and Language Processing (TSLP), 7(3):10:1– 10:??, May 2011. CODEN ???? ISSN 1550-4875.

Meinedo:2011:AGD

[MT11] Hugo Meinedo and Isabel Trancoso. Age and gender detection in the I-DASH project. ACM Transactions on Speech and Language Processing (TSLP), 7(4):13:1–13:??, August 2011. CODEN ???? ISSN 1550-4875.

Nakov:2013:SIN

[NH13] Preslav I. Nakov and Marti A.

Hearst. Semantic interpretation of noun compounds using verbal and other paraphrases. ACM Transactions on Speech and Language Processing (TSLP), 10(3):13:1–13:??, July 2013. CODEN ???? ISSN 1550-4875.

Nenkova:2007:PMI

[NPM07] Ani Nenkova, Rebecca Passonneau, and Kathleen McKeown. The Pyramid Method: Incorporating human content selection variation in summarization evaluation. ACM Transactions on Speech and Language Processing (TSLP), 4(2):4:1–4:??, May 2007. CODEN ???? ISSN 1550-4875.

Narendra:2012:SSU

[NR12] N. P. Narendra and K. Sreenivasa Rao. Syllable specific unit selection cost functions for text-to-speech synthesis. ACM Transactions on Speech and Language Processing (TSLP), 9(3):5:1–5:??, November 2012. CODEN ???? ISSN 1550-4875.

Nissim:2013:MIV

[NZ13] Malvina Nissim and Andrea Zaninello. Modeling the internal variability of multiword expressions through a pattern-based method. ACM Transactions on Speech and Language Processing (TSLP), 10(2):7:1–7:??, June 2013. CODEN ???? ISSN 1550-4875.

Oshea:2013:NBD

[OBC13] James O’shea, Zuhair Bandar, and Keeley Crockett. A new benchmark dataset with production methodology for short text semantic similarity algorithms. ACM Transactions on Speech and Language Processing (TSLP), 10(4):19:1–19:??, December 2013. CODEN ???? ISSN 1550-4875.

Pietquin:2011:SEB

[PGCFB11] Olivier Pietquin, Matthieu Geist, Senthilkumar Chandramohan, and Hervé Frezza-Buet. Sample-efficient batch reinforcement learning for dialogue management optimization. ACM Transactions on Speech and Language Processing (TSLP), 7(3):7:1–7:??, May 2011. CODEN ???? ISSN 1550-4875.

Potamianos:2011:ISI

[PGNB11] Alexandros Potamianos, Diego Giuliani, Shrikanth S. Narayanan, and Kay Berkling. Introduction to the special issue on speech and language processing of children’s speech for child-machine interaction applications. ACM Transactions on Speech and Language Processing (TSLP), 7(4):11:1–11:??, August 2011. CODEN ???? ISSN 1550-4875.

Peirsman:2011:SRB

[PP11] Yves Peirsman and Sebastian Padó. Semantic relations in bilingual lexicons. ACM Transactions on Speech and Language Processing (TSLP), 8(2): 3:1–3:??, November 2011. CO-DEN ???? ISSN 1550-4875.

Raux:2012:OTT

[RE12] Antoine Raux and Maxine Eskenazi. Optimizing the turntaking behavior of task-oriented spoken dialog systems. ACM Transactions on Speech and Language Processing (TSLP), 9 (1):1:1–1:??, May 2012. CO-DEN ???? ISSN 1550-4875.

Ringlstetter:2007:ATC

[RSM07] Christoph Ringlstetter, Klaus U. Schulz, and Stoyan Mihov. Adaptive text correction with Web-crawled domain-dependent dictionaries. ACM Transactions on Speech and Language Processing (TSLP), 4(4):9:1–9:??, October 2007. CODEN ???? ISSN 1550-4875.

Ramisch:2013:ISI

[RVK13] Carlos Ramisch, Aline Villavicencio, and Valia Kordoni. Introduction to the special issue on multiword expressions: From theory to practice and use. ACM Transactions on Speech and Language Processing (TSLP), 10(2):3:1–3:??, June 2013. CODEN ???? ISSN 1550-4875.

Rentoumi:2012:IML

[RVKM12] Vassiliki Rentoumi, George A. Vouros, Vangelis Karkaletsis, and Amalia Moser. Investigating metaphorical language in sentiment analysis: a sense-to-sentiment perspective. ACM Transactions on Speech and Language Processing (TSLP), 9(3):6:1–6:??, November 2012. CODEN ???? ISSN 1550-4875.

Roy:2013:CCN

[RZ13] Suman Deb Roy and Wenjun Zeng. Cognitive canonicalization of natural language queries using semantic strata. ACM Transactions on Speech and Language Processing (TSLP), 10(4):20:1–20:??, December 2013. CODEN ???? ISSN 1550-4875.

Shiramatsu:2008:GTM

[SKH^{+}08] Shun Shiramatsu, Kazunori Komatani, Kôiti Hasida, Tetsuya Ogata, and Hiroshi G. Okuno. A game-theoretic model of referential coherence and its empirical verification using large Japanese and English corpora. ACM Transactions on Speech and Language Processing (TSLP), 5(3):6:1–6:??, October 2008. CODEN ???? ISSN 1550-4875.

Shutova:2013:CML

[SKTK13] Ekaterina Shutova, Jakub Kaplan, Simone Teufel, and Anna Korhonen. A computational model of logical metonymy. ACM Transactions on Speech and Language Processing (TSLP), 10(3):11:1– 11:??, July 2013. CODEN ???? ISSN 1550-4875.

Sporleder:2006:BCP

[SL06] Caroline Sporleder and Mirella Lapata. Broad coverage paragraph segmentation across languages and domains. ACM Transactions on Speech and Language Processing (TSLP), 3 (2):1–35, July 2006. CODEN ???? ISSN 1550-4875.

Sakti:2012:DST

[SPF^{+}12] Sakriani Sakti, Michael Paul, Andrew Finch, Xinhui Hu, Jinfu Ni, Noriyuki Kimura, Shigeki Matsuda, Chiori Hori, Yutaka Ashikari, Hisashi Kawai, Hideki Kashioka, Eiichiro Sumita, and Satoshi Nakamura. Distributed speech translation technologies for multiparty multilingual communication. ACM Transactions on Speech and Language Processing (TSLP), 9(2):4:1– 4:??, July 2012. CODEN ???? ISSN 1550-4875.

Sokolov:2013:LBO

[SWY13] Artem Sokolov, Guillaume Wisniewski, and François Yvon. Lattice BLEU oracles in machine translation. ACM Transactions on Speech and Language Processing (TSLP), 10(4):18:1– 18:??, December 2013. CODEN ???? ISSN 1550-4875.

Editors:2013:E

[The13] The Editors. Editorial. ACM Transactions on Speech and Language Processing (TSLP), 10(4):14:1–14:??, December 2013. CODEN ???? ISSN 1550-4875.

Tomko:2005:TEH

[THT^{+}05] Stefanie Tomko, Thomas K. Harris, Arthur Toth, James Sanders, Alexander Rudnicky, and Roni Rosenfeld. Towards efficient human machine speech communication: The speech graffiti project. ACM Transactions on Speech and Language Processing (TSLP), 2(1):1–27, February 2005. CODEN ???? ISSN 1550-4875.

Turunen:2011:SRU

[TK11] Ville T. Turunen and Mikko Kurimo. Speech retrieval from unsegmented Finnish audio using statistical morpheme-like units for segmentation, recognition, and retrieval. ACM Transactions on Speech and Language Processing (TSLP), 8(1): 1:1–1:??, October 2011. CO-DEN ???? ISSN 1550-4875.

Tillmann:2007:BBP

[TZ07] Christoph Tillmann and Tong Zhang. A block bigram prediction model for statistical machine translation. ACM Transactions on Speech and Language Processing (TSLP), 4(3): 6:1–6:??, July 2007. CODEN ???? ISSN 1550-4875.

Uzeda:2010:CCE

[UPN10] Vinícius Rodrigues Uzêda, Thiago Alexandre Salgueiro Pardo, and Maria Das Graças Volpe Nunes. A comprehensive comparative evaluation of RSTbased summarization methods. ACM Transactions on Speech and Language Processing (TSLP), 6(4):4:1–4:??, May 2010. CODEN ???? ISSN 1550-4875.

VanHalteren:2007:AVL

[Van07] Hans Van Halteren. Author verification by linguistic profiling: An exploration of the parameter space. ACM Transactions on Speech and Language Processing (TSLP), 4(1):1:1–1:??, January 2007. CODEN ???? ISSN 1550-4875.

Verma:2005:VFI

[VK05] Ashish Verma and Arun Kumar. Voice fonts for individuality representation and transformation. ACM Transactions on Speech and Language Processing (TSLP), 2(1):1–19, February 2005. CODEN ???? ISSN 1550-4875.

Vincze:2013:LDE

[VTZ13] Veronika Vincze, István Nagy T., and János Zsibrita. Learning to detect English and Hungarian light verb constructions. ACM Transactions on Speech and Language Processing (TSLP), 10(2):6:1–6:??, June 2013. CODEN ???? ISSN 1550-4875.

Ward:2011:MST

[WCB^{+}11] Wayne Ward, Ronald Cole, Daniel Bolaños, Cindy Buchenroth-Martin, Edward Svirsky, Sarel Van Vuuren, Timothy Weston, Jing Zheng, and Lee Becker. My science tutor: A conversational multimedia virtual tutor for elementary school science. ACM Transactions on Speech and Language Processing (TSLP), 7(4):18:1–18:??, August 2011. CODEN ???? ISSN 1550-4875.

Wang:2006:HQS

[WS06] Chao Wang and Stephanie Seneff. High-quality speech-to-speech translation for computeraided language learning. ACM Transactions on Speech and Language Processing (TSLP), 3 (2):1–21, July 2006. CODEN ???? ISSN 1550-4875.

Wollmer:2011:TDC

[WSB^{+}11] Martin Wöllmer, Björn Schuller, Anton Batliner, Stefan Steidl, and Dino Seppi. Tandem decoding of children’s speech for keyword detection in a child-robot interaction scenario. ACM Transactions on Speech and Language Processing (TSLP), 7(4):12:1–12:??, August 2011. CODEN ???? ISSN 1550-4875.

Yan:2007:CSD

[YBKR07] Jiajun Yan, David B. Bracewell, Shingo Kuroiwa, and Fuji Ren. Chinese semantic dependency analysis: Construction of a treebank and its use in classification. ACM Transactions on Speech and Language Processing (TSLP), 4(2):5:1–5:??, May 2007. CODEN ???? ISSN 1550-4875.

Zhang:2012:ALS

[ZF12] Justin Jian Zhang and Pascale Fung. Active learning with semi-automatic annotation for extractive speech summarization. ACM Transactions on Speech and Language Processing (TSLP), 8(4):6:1–6:??, February 2012. CODEN ???? ISSN 1550-4875.

Zhang:2013:CAL

[Zha13] Li Zhang. Contextual and active learning-based affectsensing from virtual drama improvisation. ACM Transactions on Speech and Language Processing (TSLP), 9(4):8:1–8:??, January 2013. CODEN ???? ISSN 1550-4875.

Zhang:2013:TCL

[ZLG13] Renxian Zhang, Wenjie Li, and Dehong Gao. Towards contentlevel coherence with aspectguided summarization. ACM Transactions on Speech and Language Processing (TSLP), 10(1):2:1–2:??, March 2013. CODEN ???? ISSN 1550-4875.

Zhu:2012:UBA

[ZM12] Jingbo Zhu and Matthew Ma. Uncertainty-based active learning with instability estimation for text classification. ACM Transactions on Speech and Language Processing (TSLP), 8(4):5:1–5:??, February 2012. CODEN ???? ISSN 1550-4875.

Zhu:2010:CBS

[ZWHM10] Jingbo Zhu, Huizhen Wang, Eduard Hovy, and Matthew Ma. Confidence-based stopping criteria for active learning for data annotation. ACM Transactions on Speech and Language Processing (TSLP), 6(3):3:1–3:??, April 2010. CODEN ???? ISSN 1550-4875.

Zhang:2008:CWS

[ZYS08] Ruiqiang Zhang, Keiji Yasuda, and Eiichiro Sumita. Chinese word segmentation and statistical machine translation. ACM Transactions on Speech and Language Processing (TSLP), 5 REFERENCES 16

(2):4:1–4:??, May 2008. CO-DEN ???? ISSN 1550-4875.